### v2

# Deformations of pseudocharacters and Mazur’s finiteness condition

Abstract.
We show that deformation rings $R^{\mathrm{ps}}$ of $G$-pseudocharacters of a profinite group $\Gamma$ are noetherian, when $\Gamma$ satisfies Mazur’s finiteness condition. The proof proceeds by reduction to the case when $\Gamma$ is finitely generated, where the result was previously established by the second author. This enables us to extend our work on moduli spaces of $R^{\mathrm{ps}}$-condensed representations of a finitely generated profinite group $\Gamma$, to the groups satisfying Mazur’s finiteness condition. We also show that the functor from rigid analytic spaces over $\mathbb{Q}_{p}$ to sets, which associates to a rigid space $Y$ the set of continuous $\mathcal{O}(Y)$-valued $G$-pseudocharacters of $\Gamma$ is representable by a quasi-Stein rigid analytic space, and we study its general properties. We expect these results to be useful, when studying global Galois representations.
1. Introduction
If $\Omega$ is any algebraically closed field of characteristic zero then a semisimple representation $\rho:\Gamma\rightarrow\operatorname{GL}_{n}(\Omega)$ of a group $\Gamma$ is determined by its trace. Anyone who worked out a character table of a finite group can appreciate that this is a useful result. Conversely, one might wonder when a conjugation invariant function from $\Gamma$ to $\Omega$ is a trace of a representation of $\Gamma$. Lafforgue’s $G$-pseudocharacters, whose definition we recall in Section 2, provide an answer to these questions for any (generalised) reductive group scheme $G$ instead of $\operatorname{GL}_{n}$ and without restriction on the characteristic of $\Omega$. They were introduced by V. Lafforgue in his work [12] on the Langlands correspondence in the function field case in order to associate Galois representations to automorphic forms. Since then the theory has been developed further by Böckle–Harris–Khare–Thorne [3], Dat–Helm–Kurinczuk–Moss [8], Emerson–Morel [10], Weidner [21], Zhu [22], and J.Q. [16].
1.1. Results
Let $L$ be a finite extension of $\mathbb{Q}_{p}$ with a ring of integers $\mathcal{O}$ and residue field $k$. Let $G$ be a (generalised) reductive group scheme over $\mathcal{O}$, let $G^{0}$ be the neutral component of $G$. Let $\Gamma$ be a profinite group and let $\overline{\Theta}$ be a continuous $k$-valued $G$-pseudocharacter. It is shown in [16] that the deformation problem $\operatorname{Def}^{\Gamma}_{\overline{\Theta}}$, parameterising continuous $A$-valued $G$-pseudocharacters of $\Gamma$ lifting $\overline{\Theta}$, where $A$ is a local artinian $\mathcal{O}$-algebra with residue field $k$, is pro-represented by a local profinite $\mathcal{O}$-algebra $R^{\mathrm{ps},\Gamma}_{\overline{\Theta}}$ with residue field $k$.
By the reconstruction theorem alluded to above there is a continuous representation $\overline{\rho}:\Gamma\rightarrow G(\bar{k})$, such that the $G$-pseudocharacter $\Theta_{\overline{\rho}}$ associated to $\overline{\rho}$ is equal to $\overline{\Theta}$. Let $K$ be the kernel of $\overline{\rho}$ and let $Q$ be a quotient of $\Gamma$ fitting into an exact sequence
(1) $0\rightarrow K(p)\rightarrow Q\rightarrow\Gamma/K\rightarrow 0,$
where $K(p)$ is the maximal pro-$p$ quotient of $K$. We show in 5.3 that $Q$ does not depend on the choice of $\overline{\rho}$. We may view $\overline{\Theta}$ as a $G$-pseudocharacter of $Q$. Moreover, any continuous pseudocharacter of $Q$ becomes a continuous pseudocharacter of $\Gamma$ after composing it with the quotient map. Hence, $\operatorname{Def}^{Q}_{\overline{\Theta}}$ is a closed subfunctor of $\operatorname{Def}^{\Gamma}_{\overline{\Theta}}$, which induces a surjection $R^{\mathrm{ps},\Gamma}_{\overline{\Theta}}\twoheadrightarrow R^{\mathrm{ps},Q}_{\overline{\Theta}}$. The following theorem is the technical heart of the paper.
Theorem 1.1 (5.6).
The map $R^{\mathrm{ps},\Gamma}_{\overline{\Theta}}\twoheadrightarrow R^{\mathrm{ps},Q}_{\overline{\Theta}}$ is an isomorphism. In particular, every deformation of $\overline{\Theta}$ factors through $Q$.
A profinite group $\Gamma$ satisfies Mazur’s $p$-finiteness condition $\Phi_{p}$ if for every open subgroup $\Gamma^{\prime}$ of $\Gamma$ the space of continuous group homomorphisms $\operatorname{Hom}^{\mathrm{cont}}(\Gamma^{\prime},\mathbb{F}_{p})$ is finite. If $\Gamma$ satisfies $\Phi_{p}$ then it is easy to see that $K(p)$ is a finitely generated pro-$p$ subgroup and hence $Q$ is a finitely generated [^1] Throughout the paper this means that the group in question is finitely generated as a topological group. profinite group. It was shown in [16] that the ring $R^{\mathrm{ps},Q}_{\overline{\Theta}}$ is noetherian and hence we obtain:
Corollary 1.2.
If $\Gamma$ satisfies Mazur’s condition $\Phi_{p}$ then $R^{\mathrm{ps},\Gamma}_{\overline{\Theta}}$ is noetherian.
In [14] for a finitely generated profinite group $\Gamma$ we have defined a functor $X^{\mathrm{gen},\Gamma}_{\overline{\Theta}}:R^{\mathrm{ps},\Gamma}_{\overline{\Theta}}\text{-}\mathrm{alg}\rightarrow\operatorname{Set}$, such that $X^{\mathrm{gen},\Gamma}_{\overline{\Theta}}(A)$ is the set of $R^{\mathrm{ps}}_{\overline{\Theta}}$-condensed representations $\rho:\Gamma\rightarrow G(A)$, such that the $G$-pseudocharacter $\Theta_{\rho}$ of $\rho$ is equal to the universal deformation $\Theta^{u}$ specialised along the map $R^{\mathrm{ps},\Gamma}_{\overline{\Theta}}\rightarrow A$. The condition $R^{\mathrm{ps},\Gamma}_{\overline{\Theta}}$-condensed is a continuity condition on the representation expressed in an algebraic way. We have shown in [14] that for finitely generated profinite groups $\Gamma$ the functor $X^{\mathrm{gen},\Gamma}_{\overline{\Theta}}$ is represented by a finite type $R^{\mathrm{ps}}_{\overline{\Theta}}$-algebra $A^{\mathrm{gen},\Gamma}_{\overline{\Theta}}$. 1.2 allows us to define $X^{\mathrm{gen},\Gamma}_{\overline{\Theta}}$ for profinite groups $\Gamma$ satisfying $\Phi_{p}$ in the same way as for finitely generated profinite groups. Using 1.1 we prove:
Theorem 1.3 (6.3).
If $\Gamma$ satisfies $\Phi_{p}$, then $X^{\mathrm{gen},\Gamma}_{\overline{\Theta}}=X^{\mathrm{gen},Q}_{\overline{\Theta}}$.
Since $Q$ is finitely generated the above theorem allows us to transfer the results concerning $X^{\mathrm{gen},\Gamma}_{\overline{\Theta}}$ proved for finitely generated profinite groups $\Gamma$ in [14], [15] to the profinite groups satisfying $\Phi_{p}$. In particular, we obtain
Corollary 1.4.
If $\Gamma$ satisfies Mazur’s condition $\Phi_{p}$, then $X^{\mathrm{gen},\Gamma}_{\overline{\Theta}}$ is represented by a finite type $R^{\mathrm{ps},\Gamma}_{\overline{\Theta}}$-algebra.
To a representation $\rho:\Gamma\rightarrow G(A)$ one may associate its $G$-pseudocharacter $\Theta_{\rho}$. This map induces a morphism of schemes $X^{\mathrm{gen},\Gamma}_{\overline{\Theta}}\rightarrow X^{\mathrm{ps},\Gamma}_{\overline{\Theta}}$, where $X^{\mathrm{ps},\Gamma}_{\overline{\Theta}}=\operatorname{Spec}R^{\mathrm{ps},\Gamma}_{\overline{\Theta}}$. The morphism is $G^{0}$-equivariant for the conjugation action on the source and the trivial action on the target. Thus it factors through the GIT-quotient $X^{\mathrm{gen},\Gamma}_{\overline{\Theta}}\sslash G^{0}\rightarrow X^{\mathrm{ps}}_{\overline{\Theta}}$. By appealing to the analogous result proved for finitely generated groups $\Gamma$ in [15] we obtain:
Corollary 1.5.
If $\Gamma$ satisfies $\Phi_{p}$, then $X^{\mathrm{gen},\Gamma}_{\overline{\Theta}}\sslash G^{0}\rightarrow X^{\mathrm{ps}}_{\overline{\Theta}}$ is a finite adequate homeomorphism in the sense of [1, Definition 3.3.1].
If $G=\operatorname{GL}_{n}$ this result was proved by Carl Wang-Erickson [20] building on the results of Gaëtan Chenevier on determinant laws [6].
Let $\mathrm{Rig}_{L}$ be the category of rigid analytic spaces over $L$. If $Y\in\mathrm{Rig}_{L}$ then its ring of global sections $\mathcal{O}(Y)$ carries a natural topology. We let $\tilde{X}_{G}:\mathrm{Rig}_{L}^{\mathrm{op}}\rightarrow\operatorname{Set}$ be the functor which to a rigid analytic space $Y$ associates the set of continuous $G$-pseudocharacters $\Theta$ valued in $\mathcal{O}(Y)$.
Theorem 1.6 (8.3, 8.4).
If $\Gamma$ satisfies Mazur’s condition $\Phi_{p}$, then the functor $\tilde{X}_{G}$ is representable by a quasi-Stein rigid analytic space $X_{G}$. The $\overline{L}$-points of $X_{G}$ are in canonical bijection with $G^{0}(\overline{L})$-conjugacy classes of continuous $G$-semisimple [^2] We remind the reader that a representation $\rho:\Gamma\rightarrow G(\Omega)$, where $\Omega$ is an algebraically closed field, is $G$-semisimple (or $G$-completely reducible) provided that whenever $\rho(\Gamma)$ is contained in $P(\Omega)$ for a parabolic subgroup $P$ of $G$, it is contained in $L(\Omega)$ for a Levi subgroup $L$ of $P$. We refer the reader to [2, Section 6] for more details. representations $\rho:\Gamma\rightarrow G(\overline{L})$.
If $\Gamma$ is finitely generated then the result was proved by J.Q. in [16]. Our proof proceeds by reduction to the finitely generated case. To carry out the reduction in 7.3 we prove an analog of 1.1 for deformation rings of continuous $G$-pseudocharacters valued in local fields. If $G=\operatorname{GL}_{n}$ then 1.6 is the main result of [6].
1.2. What was known and tried before
Let us recall the previous related results. If $G=\operatorname{GL}_{n}$ then using the work of Emerson–Morel [10] one can work with Chenevier’s determinant laws instead of $\operatorname{GL}_{n}$-pseudocharacters. In this context 1.1 is proven by Chenevier in [6]: as part of the theory one gets a Cayley–Hamilton algebra $E$ together with a continuous representation $\rho^{\mathrm{univ}}:\Gamma\rightarrow E^{\times}$. One may then show that $\rho^{\mathrm{univ}}$ factors through $Q$ and deduce 1.1 from this, see [6, Lemma 3.8] and [14, Lemma 5.5] for details.
If $G$ is any generalised reductive group then the Cayley–Hamilton algebra is not available anymore. However, one could still try to deduce 1.2 from the $\operatorname{GL}_{n}$-case by choosing a closed immersion $\tau:G\hookrightarrow\operatorname{GL}_{n}$ and studying the map $R^{\mathrm{ps},\Gamma}_{\tau\circ\overline{\Theta}}\rightarrow R^{\mathrm{ps},\Gamma}_{\overline{\Theta}}$ induced by composition with $\tau$. If the map is finite then one may deduce from the $\operatorname{GL}_{n}$-case that $R^{\mathrm{ps},\Gamma}_{\overline{\Theta}}$ is noetherian. This approach was carried out by J.Q. in [16, Section 5.3] for classical groups under the assumption $p>2$ in the orthogonal cases. However, this strategy runs into difficult invariant theoretic questions in general. Since $R^{\mathrm{ps},\Gamma}_{\tau\circ\overline{\Theta}}\rightarrow R^{\mathrm{ps},\Gamma}_{\overline{\Theta}}$ is in general not surjective the authors do not see a way of deducing 1.1 from the $\operatorname{GL}_{n}$-case.
If $\Gamma$ is a finitely generated profinite group then J.Q. proves 1.2 by showing that the tangent space is finite dimensional, which is equivalent to showing that $\operatorname{Def}_{\overline{\Theta}}^{\Gamma}(k[\varepsilon])$ is finite. Let us informally explain the underlying difficulty if we drop the assumption that $\Gamma$ is finitely generated. The $G$-pseudocharacter $\Theta$ valued in the ring of dual numbers $k[\varepsilon]$ over $k$ consists of $\mathcal{O}$-algebra homomorphisms $\Theta_{n}:\mathcal{O}[G^{n}]^{G^{0}}\rightarrow\mathcal{C}(\Gamma^{n},k[\varepsilon])$ for $n\geq 1$ satisfying certain compatibility relations, where $\mathcal{C}(\Gamma^{n},k[\varepsilon])$ is the set of continuous maps. Proving finiteness assertions such as finite dimensionality of the tangent space amounts to showing that finitely many $\Theta_{n}$ determine $\Theta$ uniquely. This is feasible when the group is finitely generated, but the argument runs into trouble when this assumption is dropped. A similar problem arises in 1.6. One has to show that for an affinoid algebra $A$ a continuous $A$-valued $G$-pseudocharacter $\Theta$ takes values in some formal model of $A$. It is easy to show that each $\Theta_{n}$ takes values in $\mathcal{C}(\Gamma^{n},\mathcal{A}_{n})$ for some formal model $\mathcal{A}_{n}$ of $A$, but it is not clear that there is a formal model that works for all $\Theta_{n}$.
1.3. What we actually do
Let us explain the proof of 1.1. We first prove 1.1 for finitely generated free profinite groups $\Gamma$. We show that the map between deformation rings $R^{\mathrm{ps}}_{\overline{\Theta}}\rightarrow R^{\square}_{\overline{\rho}}$, induced by mapping a deformation of $\overline{\rho}$ to its $G$-pseudocharacter, is injective. This key step is carried out in Section 4. Its proof uses the fact that one may describe the moduli space of $G$-pseudocharacters of a discrete finitely generated free group $F$ as a GIT-quotient of a moduli space of representations of $F$ into $G$. We show that if $\Gamma$ is the profinite completion of $F$ then the rings $R^{\mathrm{ps}}_{\overline{\Theta}}$ and $R^{\square}_{\overline{\rho}}$ are naturally isomorphic to completions of the local rings at points $\Theta|_{F}$ and $\overline{\rho}|_{F}$ in the respective moduli spaces. The assertion about injectivity boils down to commutative algebra, which is carried out in Section 3. We note that if we drop the assumption that $\Gamma$ is a free profinite group then it is not known [^3] Carl Wang-Erickson has pointed out to us that Jinyue Luo [13] has constructed an example, where the map fails to be injective. whether $R^{\mathrm{ps}}_{\overline{\Theta}}\rightarrow R^{\square}_{\overline{\rho}}$ is injective, even when $G=\operatorname{GL}_{n}$, see [6, Remark 1.25].
It is easy to see that the universal deformation $\rho^{\mathrm{univ}}:\Gamma\rightarrow G(R^{\square}_{\overline{\rho}})$ factors through $Q$ and the injectivity of the map $R^{\mathrm{ps}}_{\overline{\Theta}}\rightarrow R^{\square}_{\overline{\rho}}$ implies 1.1 for finitely generated free profinite groups $\Gamma$. From this one may deduce 1.1 for all finitely generated profinite groups $\Gamma$.
If $\Theta$ is a continuous $G$-pseudocharacter of $\Gamma$ valued in a Hausdorff topological ring $A$ then there is a closed normal subgroup $\operatorname{Ker}(\Theta)$ of $\Gamma$, whose definition we recall in Section 2.2, such that $\Theta$ factors through $\Gamma/\operatorname{Ker}(\Theta)$. 1.1 holds if and only if $K/\operatorname{Ker}(\Theta^{u})$ is a pro-$p$ group, where $\Theta^{u}$ is the universal deformation of $\overline{\Theta}$. A profinite group is pro-$p$ if and only if every closed finitely generated subgroup is pro-$p$. This allows us to prove that $K/\operatorname{Ker}(\Theta^{u})$ is pro-$p$ for general $\Gamma$ by considering the restrictions of $\Theta^{u}$ to closed finitely generated subgroups of $\Gamma$ and using the fact that we have already proved 1.1 for such groups.
1.4. Motivation
The Hermite–Minkowski theorem implies that the Galois group $\operatorname{Gal}(F_{S}/F)$, where $F$ is a number field and $F_{S}$ is a maximal extension of $F$ unramified outside a finite set of primes $S$ of $F$, satisfies $\Phi_{p}$. It is not known if these groups are topologically finitely generated. The deformation rings of representations of $\operatorname{Gal}(F_{S}/F)$ are frequently encountered in arithmetic applications and $\Gamma=\operatorname{Gal}(F_{S}/F)$ is the main example that we have in mind in this paper.
In [14] we used the moduli space $X^{\mathrm{gen},\Gamma}_{\overline{\Theta}}$ to study the deformation rings $R^{\square}_{\rho}$ of representations $\rho:\Gamma\rightarrow G(k)$ such that $\Theta_{\rho}=\overline{\Theta}$, when $\Gamma$ is absolute Galois group of a $p$-adic local field. We expect the results of this paper to be useful, when studying analogous questions for $\Gamma=\operatorname{Gal}(F_{S}/F)$ for a number field $F$.
1.5. Acknowledgements
V.P. thanks Ulrich Görtz and Marc Levine for a stimulating discussion regarding Section 3. The authors thank Gebhard Böckle, Toby Gee and Carl Wang-Erickson for their comments and the referee for their careful reading of the paper.
The research of J.Q. was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – project number 517234220.
2. Lafforgue’s pseudocharacters
In this section we recall the definition of pseudocharacters and their deformation theory. Until Section 2.3 we let $\mathcal{O}$ be any commutative ring.
A generalised reductive group $G$ over $\mathcal{O}$ is a smooth affine group scheme, that has reductive geometric fibres and such that $G/G^{0}$ is finite.
Definition 2.1.
Let $\Gamma$ be an abstract group and let $A$ be a commutative $\mathcal{O}$-algebra. An $G$-pseudocharacter $\Theta$ of $\Gamma$ over $A$ is a sequence $(\Theta_{n})_{n\geq 1}$ of $\mathcal{O}$-algebra maps
$\Theta_{n}:\mathcal{O}[G^{n}]^{G^{0}}\to\mathrm{Map}(\Gamma^{n},A)$
for $n\geq 1$, satisfying the following conditions [^4] Here $G$ acts on $G^{n}$ by $g\cdot(g_{1},\dots,g_{n})=(gg_{1}g^{-1},\dots,gg_{n}g^{-1})$. This induces a rational action of $G$ on the affine coordinate ring $\mathcal{O}[G^{n}]$ of $G^{n}$. The submodule $\mathcal{O}[G^{n}]^{G^{0}}\subseteq\mathcal{O}[G^{n}]$ is defined as the algebraic invariant module of the $G^{0}$-representation $\mathcal{O}[G^{n}]$. It is an $\mathcal{O}$-subalgebra, since $G$ acts by $\mathcal{O}$-linear automorphisms.:
  1. (1)
    For each $n,m\geq 1$, each map $\zeta:\{1,\dots,m\}\to\{1,\dots,n\}$, $f\in\mathcal{O}[G^{m}]^{G^{0}}$ and $\gamma_{1},\dots,\gamma_{n}\in\Gamma$, we have
    $\Theta_{n}(f^{\zeta})(\gamma_{1},\dots,\gamma_{n})=\Theta_{m}(f)(\gamma_{\zeta(1)},\dots,\gamma_{\zeta(m)})$
    where $f^{\zeta}(g_{1},\dots,g_{n})=f(g_{\zeta(1)},\dots,g_{\zeta(m)})$.
  2. (2)
    For each $n\geq 1$, for each $\gamma_{1},\dots,\gamma_{n+1}\in\Gamma$ and each $f\in\mathcal{O}[G^{n}]^{G^{0}}$, we have
    $\Theta_{n+1}(\hat{f})(\gamma_{1},\dots,\gamma_{n+1})=\Theta_{n}(f)(\gamma_{1},\dots,\gamma_{n-1},\gamma_{n}\gamma_{n+1})$
    where $\hat{f}(g_{1},\dots,g_{n+1})=f(g_{1},\dots,g_{n-1},g_{n}g_{n+1})$.
We denote the set of $G$-pseudocharacters of $\Gamma$ over $A$ by $\mathrm{PC}_{G}^{\Gamma}(A)$. If $f:A\to B$ is a homomorphism of $\mathcal{O}$-algebras, then there is an induced map $f_{*}:\mathrm{PC}_{G}^{\Gamma}(A)\to\mathrm{PC}_{G}^{\Gamma}(B)$. For $\Theta\in\mathrm{PC}_{G}^{\Gamma}(A)$, the image $f_{*}(\Theta)$ is called the specialisation of $\Theta$ along $f$ and is denoted by $\Theta\otimes_{A}B$. It is easy to verify that specialisation along $f:A\to B$ commutes with composition with $\varphi$, i.e. $(\varphi\circ\Theta)\otimes_{A}B=\varphi\circ(\Theta\otimes_{A}B)$.
The functor $A\mapsto\mathrm{PC}^{\Gamma}_{G}(A)$ is representable by an $\mathcal{O}$-algebra [16, Theorem 3.19]. Let us write $\operatorname{Rep}^{\Gamma}_{G}(A):=\operatorname{Hom}(\Gamma,G(A))$. The functor $A\mapsto\operatorname{Rep}^{\Gamma}_{G}(A)$ is also representable by an $\mathcal{O}$-algebra. To a representation $\rho:\Gamma\to G(A)$, we associate a $G$-pseudocharacter $\Theta_{\rho}\in\mathrm{PC}^{\Gamma}_{G}(A)$ by the formula
(2) $\Theta_{\rho,m}(f)(\gamma_{1},\dots,\gamma_{m}):=f(\rho(\gamma_{1}),\dots,\rho(\gamma_{m}))$
for all $m\geq 1$, all $f\in\mathcal{O}[G^{m}]^{G^{0}}$ and all $\gamma_{1},\dots,\gamma_{m}\in\Gamma$. The map $\rho\mapsto\Theta_{\rho}$ defines a morphism of schemes $\operatorname{Rep}^{\Gamma}_{G}\to\mathrm{PC}^{\Gamma}_{G}$, which is $G^{0}$-equivariant for the conjugation action on the source and the trivial action on the target and hence induces a morphism from the GIT-quotient $\operatorname{Rep}^{\Gamma}_{G}\sslash G^{0}\rightarrow\mathrm{PC}^{\Gamma}_{G}$.
2.1. Continuous pseudocharacters
If $\Gamma$ is a topological group, $A$ is a topological ring and, for all $n\geq 1$ the map $\Theta_{n}$ has image in the set $\mathcal{C}(\Gamma^{n},A)$ of continuous maps $\Gamma^{n}\to A$, then we say that $\Theta$ is continuous. We denote the set of continuous $G$-pseudocharacters of $\Gamma$ with values in $A$ by $\mathrm{cPC}_{G}^{\Gamma}(A)$.
2.2. Kernels
If $\Theta\in\mathrm{PC}^{\Gamma}_{G}(A)$ then the kernel $\operatorname{Ker}(\Theta)$ of $\Theta$ is defined as the set of all $\delta\in\Gamma$, such that for all $n\geq 1$, all $f\in\mathcal{O}[G^{n}]^{G^{0}}$ and all $\gamma_{1},\ldots,\gamma_{n}\in\Gamma$, we have
(3) $\Theta_{n}(f)(\gamma_{1},\ldots,\gamma_{n}\delta)=\Theta_{n}(f)(\gamma_{1},\ldots,\gamma_{n}).$
It is shown in [16, Lemma 3.11] that $\operatorname{Ker}(\Theta)$ is a normal subgroup of $\Gamma$.
Lemma 2.2.
Assume that $\Gamma$ is a topological group, $A$ is a topological ring and $\Theta\in\mathrm{cPC}^{\Gamma}_{G}(A)$. If $A$ is Hausdorff (e.g. profinite) then $\operatorname{Ker}(\Theta)$ is closed in $\Gamma$.
Proof.
For fixed $(\gamma_{1},\ldots,\gamma_{n})\in\Gamma^{n}$ and $f\in\mathcal{O}[G^{n}]^{G^{0}}$ the map
(4) $\Gamma\rightarrow A,\quad\delta\mapsto\Theta_{n}(f)(\gamma_{1},\ldots,\gamma_{n}\delta)-\Theta_{n}(f)(\gamma_{1},\ldots,\gamma_{n})$
is continuous. If $A$ is Hausdorff then the preimage of $\{0\}$ is a closed subset of $\Gamma$. Since $\operatorname{Ker}(\Theta)$ is the intersection of all such subsets we obtain the assertion. ∎
2.3. Deformations of $G$-pseudocharacters
Let $\mathcal{O}$ be the ring of integers in a finite extension $L$ of $\mathbb{Q}_{p}$ with residue field $k$. Let $\Gamma$ be a profinite group and let $\overline{\Theta}\in\mathrm{cPC}^{\Gamma}_{G}(k)$. We recall a deformation problem of $\overline{\Theta}$ introduced in [16, Section 5]. Let $\mathfrak{A}_{\mathcal{O}}$ be the category of artinian local $\mathcal{O}$-algebras with residue field $k$. We define the deformation functor of $\overline{\Theta}$
$\operatorname{Def}_{\overline{\Theta}}:\mathfrak{A}_{\mathcal{O}}\to\operatorname{Set},~A\mapsto\{\Theta\in\mathrm{cPC}_{G}^{\Gamma}(A)\mid\Theta\otimes_{A}k=\overline{\Theta}\}$
that sends an object $A\in\mathfrak{A}_{\mathcal{O}}$ to the set of continuous $G$-pseudocharacters $\Theta$ of $\Gamma$ over $A$ with $\Theta\otimes_{A}k=\overline{\Theta}$. It is pro-representable by a local profinite $\mathcal{O}$-algebra $R^{\mathrm{ps}}_{\overline{\Theta}}$ with residue field $k$ by [16, Theorem 5.4].
We will add a superscript $\Gamma$ and write $\operatorname{Def}^{\Gamma}_{\overline{\Theta}}$, $R^{\mathrm{ps},\Gamma}_{\overline{\Theta}}$ to emphasise that we are working with pseudocharacters of $\Gamma$, and will drop the superscript if the context is clear.
Denote by $\Theta^{u}\in\mathrm{cPC}^{\Gamma}_{G}(R^{\mathrm{ps}}_{\overline{\Theta}})$ the universal deformation of $\overline{\Theta}$. If $A$ is an $R^{\mathrm{ps}}_{\overline{\Theta}}$-algebra we will write $\Theta^{u}_{|A}$ for the specialisation of $\Theta^{u}$ along $R^{\mathrm{ps}}_{\overline{\Theta}}\rightarrow A$. Since $R^{\mathrm{ps}}_{\overline{\Theta}}$ is profinite and hence Hausdorff, $\operatorname{Ker}(\Theta^{u})$ is a closed normal subgroup of $\Gamma$ by 2.2.
3. Geometric invariant theory
We first recall the set up of [18]. Let $\mathcal{O}$ be a quasi-excellent ring and let $S=\operatorname{Spec}\mathcal{O}$. Then $\mathcal{O}$ is Nagata [19, Tag 07QV] and hence universally Japanese [19, Tag 0334]. The last assumption is imposed in [18]. Let $G$ be a reductive group scheme over $S$, so that $G$ is an affine group scheme over $S$, $G\rightarrow S$ is smooth and the geometric fibres are connected reductive groups.
Let $V$ be a free $\mathcal{O}$-module of finite rank $r$ endowed with a $G$-module structure, let $\check{V}=\operatorname{Hom}_{\mathcal{O}}(V,\mathcal{O})$ and let $\operatorname{Sym}(\check{V})$ be the symmetric algebra over $\mathcal{O}$. The $G$-module structure on $V$ induces an action of $G$ on $\operatorname{Spec}(\operatorname{Sym}(\check{V}))=\mathbb{A}^{r}_{S}.$ Let $X$ be a closed $G$-invariant subscheme of $\operatorname{Spec}(\operatorname{Sym}(\check{V}))$. The $G$-action on $X$ induces an action on $B$, the ring of functions on $X$. The GIT quotient $X\sslash G$ is represented by the ring of invariants $B^{G}$. Below we let $R$ be a $B^{G}$-algebra, which we assume to be a quasi-excellent ring. In the application, both $\mathcal{O}$ and $R$ will be complete noetherian local rings, and these are quasi-excellent [19, Tag 07QW].
Lemma 3.1.
If $R$ is a flat $B^{G}$-algebra or a flat $\mathcal{O}$-algebra then the ring $(R\otimes_{B^{G}}B)^{G}$ is a finite $R$-module.
Proof.
If $R$ is a flat $B^{G}$-algebra then [18, Lemma 2] implies that
(5) $(R\otimes_{B^{G}}B)^{G}\cong R\otimes_{B^{G}}B^{G}\cong R$
and we are done. If $R$ is flat over $\mathcal{O}$ then [18, Lemma 2] implies that
(6) $(R\otimes_{\mathcal{O}}\operatorname{Sym}(\check{V}))^{G}\cong R\otimes_{\mathcal{O}}\operatorname{Sym}(\check{V})^{G}.$
It follows from [18, Theorem 2 (ii)] that $(R\otimes_{B^{G}}B)^{G}$ is a finite $(R\otimes_{\mathcal{O}}\operatorname{Sym}(\check{V}))^{G}$-module. Since the image of $\operatorname{Sym}(\check{V})^{G}$ in $B$ is contained in $B^{G}$, the action of $R\otimes_{\mathcal{O}}\operatorname{Sym}(\check{V})^{G}$ on $(R\otimes_{B^{G}}B)^{G}$ factors through the map $R\otimes_{\mathcal{O}}\operatorname{Sym}(\check{V})^{G}\rightarrow R\otimes_{\mathcal{O}}B^{G}\rightarrow R$. Hence, $(R\otimes_{B^{G}}B)^{G}$ is a finite $R$-module. ∎
Lemma 3.2.
Let $R$ be a local ring with residue field $\kappa$. Assume that $R$ is either $B^{G}$-flat or $\mathcal{O}$-flat. Let $Z$ be a closed non-empty subscheme of $X\times_{X\sslash G}\operatorname{Spec}R$. If $Z$ is $G$-invariant then every irreducible component of $Z$ contains a point $x$ above the closed point of $\operatorname{Spec}R$ such that $\kappa(x)$ is finite over $\kappa$.
Proof.
3.1 implies $(R\otimes_{B^{G}}B)^{G}$ is a finite $R$-module. Since $R$ is a local noetherian ring this implies that every closed subscheme of $(X\times_{X\sslash G}\operatorname{Spec}R)\sslash G$ contains a point above the closed point of $\operatorname{Spec}R$. Since $G$ is connected every irreducible component of $Z$ is $G$-invariant by [4, Lemma 2.1]. Thus we may assume that $Z$ is irreducible. The image of $Z$ in $(X\times_{X\sslash G}\operatorname{Spec}R)\sslash G$ is closed by [18, Theorem 3 (iii)] and thus $Z$ contains a point $x$ above the closed point of $\operatorname{Spec}R$. Since $X\rightarrow X\sslash G$ is of finite type we may choose $x$ such that $\kappa(x)$ is finite over $\kappa$. ∎
Proposition 3.3.
Let $\bar{x}:\operatorname{Spec}\Omega\rightarrow X$ be a geometric point, such that its $G$-orbit in $X\otimes_{\mathcal{O}}\Omega$ is closed. Let $x$ be its image in $X$ and let $y$ be its image in $X\sslash G$. If $X$ is normal then the induced map on the completions of local rings $\widehat{\mathcal{O}}_{X\sslash G,y}\rightarrow\widehat{\mathcal{O}}_{X,x}$ is injective.
Proof.
Let $\mathfrak{p}$ be the prime ideal of $B^{G}$ corresponding to $y$ and let $R$ be the completion of the localisation $(B^{G})_{\mathfrak{p}}$. We note that $R$ is complete local noetherian and hence quasi-excellent and also $B^{G}$-flat. Since $B$ is normal, $B^{G}$ and $(B^{G})_{\mathfrak{p}}$ are normal. Since $\mathcal{O}$ is quasi-excellent and $B^{G}$ is of finite type over $\mathcal{O}$ by [18, Theorem 2], $(B^{G})_{\mathfrak{p}}$ is quasi-excellent [19, Tag 07QU]. Since $(B^{G})_{\mathfrak{p}}$ is quasi-excellent and normal, its completion $R$ is also normal [19, Tag 0C23], and hence an integral domain.
Let $Q$ be the fraction field of $R$. Since $R$ is $B^{G}$-flat and $Q$ is $R$-flat, $R$ is a subring of $B_{R}:=R\otimes_{B^{G}}B$ and $Q$ is a subring of $B_{Q}:=Q\otimes_{B^{G}}B$. In particular, $X_{R}:=\operatorname{Spec}B_{R}$ and $X_{Q}:=\operatorname{Spec}B_{Q}$ are non-empty. Let $Z^{\prime}$ be an irreducible component of $X_{Q}$ corresponding to a prime ideal $\mathfrak{q}^{\prime}$ of $B_{Q}$. Then $Z^{\prime}$ is $G$-invariant by [4, Lemma 2.1]. Let $Z$ be the closure of $Z^{\prime}$ in $X_{R}$. Then $Z$ is a closed $G$-invariant subset of $X_{R}$, corresponding to the prime ideal $\mathfrak{q}^{\prime\prime}=\mathfrak{q}^{\prime}\cap B_{R}$.
3.2 implies that there is $\bar{x}_{1}\in Z(\Omega)$ above $y$. Let $\bar{x}_{2}\in Z(\Omega)$ be a point contained in the closure of the $G$-orbit of $\bar{x}_{1}$ in $Z\otimes_{R}\Omega$, such that its $G$-orbit has minimal dimension. Then $\bar{x}_{2}\in Z(\Omega)$ lies above $y$ and its $G$-orbit is closed. It follows from [18, Theorem 3 (ii)] that $\bar{x}$ and $\bar{x}_{2}$ lie in the same $G$-orbit and thus $\bar{x}\in Z(\Omega)$.
Let $\mathfrak{n}$ be a prime ideal of $B_{R}$ corresponding to the image of $\bar{x}$ in $X_{R}$ and let $\mathfrak{q}$ be the prime ideal of $B$ corresponding to $x$. We have to show that the natural map $R\rightarrow\widehat{B}_{\mathfrak{q}}$ is injective. Since $R$ is the completion of $(B^{G})_{\mathfrak{p}}$ the completions of $(R\otimes_{B^{G}}B)_{\mathfrak{n}}$ and $B_{\mathfrak{q}}$ coincide. Thus it is enough to show that the map $R\rightarrow(B_{R})_{\mathfrak{n}}$ is injective.
By construction we have inclusions $R\subset Q\subset B_{Q}/\mathfrak{q}^{\prime}\subset\kappa(\mathfrak{q}^{\prime})$. Thus the composition $R\rightarrow(B_{R})_{\mathfrak{n}}\rightarrow(B_{R}/\mathfrak{q}^{\prime\prime})_{\mathfrak{n}}\subset\kappa(\mathfrak{q}^{\prime})$ is injective, which implies that $R\rightarrow(B_{R})_{\mathfrak{n}}$ is injective. ∎
4. Finitely generated free profinite groups
Let $F$ be a free group on $r$ generators $\gamma_{1},\ldots,\gamma_{r}$. Let $L$ be a finite extension of $\mathbb{Q}_{p}$ with the ring of integers $\mathcal{O}$ and residue field $k$. Let $G$ be a generalised reductive group scheme over $\mathcal{O}$.
Proposition 4.1.
The natural map $\operatorname{Rep}^{F}_{G}\sslash G^{0}\rightarrow\mathrm{PC}^{F}_{G}$ is an isomorphism.
Proof.
By [16, Theorem 3.20], $\mathrm{PC}^{F}_{G}$ is representable by $\mathop{\mathstrut\rm colim}\limits\nolimits_{F_{m}\to F}\mathcal{O}[G^{m}]^{G^{0}}$, where the colimit is indexed by the category of finitely generated free groups $F_{m}$ together with a homomorphism $F_{m}\to F$ and homomorphisms of free groups commuting with the homomorphisms to $F$. This category has a terminal object, given by an isomorphism $F_{r}\xrightarrow{\cong}F$. It follows, that the colimit is isomorphic to $\mathcal{O}[G^{r}]^{G^{0}}$. ∎
Let $\overline{\Theta}\in\mathrm{PC}^{F}_{G}(k)$ and let $\overline{\rho}:F\rightarrow G(\bar{k})$ be a $G$-semisimple representation such that $\Theta_{\overline{\rho}}=\overline{\Theta}\otimes_{k}\bar{k}$, such representation exists and is uniquely determined up to $G^{0}(\bar{k})$-conjugation by [16, Theorem 3.7].
Proposition 4.2.
Let $\mathfrak{m}$ be a maximal ideal of $\mathcal{O}(\mathrm{PC}^{F}_{G})$ corresponding to $\overline{\Theta}$ and let $\mathfrak{n}$ be a maximal ideal of $\mathcal{O}(\operatorname{Rep}^{F}_{G})$ corresponding to $\overline{\rho}$. Then the morphism in 4.1 induces an injection on completions of local rings $\mathcal{O}(\mathrm{PC}^{F}_{G})_{\mathfrak{m}}^{\wedge}\hookrightarrow\mathcal{O}(\operatorname{Rep}^{F}_{G})^{\wedge}_{\mathfrak{n}}$.
Proof.
The map $\rho\mapsto(\rho(\gamma_{1}),\ldots,\rho(\gamma_{r}))$ induces an isomorphism of $\mathcal{O}$-schemes $\operatorname{Rep}^{F}_{G}\cong G^{r}$. Since $G$ is smooth over $\mathcal{O}$, we deduce that $\operatorname{Rep}^{F}_{G}$ is smooth over $\mathcal{O}$. Since $\mathcal{O}$ is a DVR we deduce that $\operatorname{Rep}^{F}_{G}$ is regular and hence normal. Since $\overline{\rho}$ is $G$-semisimple the $G$-orbit of $\bar{x}\in\operatorname{Rep}^{F}_{G}(\bar{k})$ corresponding to $\overline{\rho}$ is closed, [14, Proposition 6.7]. The assertion follows from 4.1 and 3.3 applied with $X=\operatorname{Rep}^{F}_{G}$. ∎
There is a finite extension $k^{\prime}$ of $k$ inside $\bar{k}$ such that all $\overline{\rho}(\gamma_{i})$ are contained in $G(k^{\prime})$. Then $\overline{\rho}(F)\subset G(k^{\prime})$ and since $k^{\prime}$ is a finite field the homomorphism $\overline{\rho}:F\rightarrow G(\bar{k})$ extends to a continuous homomorphism $\overline{\rho}:\widehat{F}\rightarrow G(\bar{k})$, where $\widehat{F}$ is the profinite completion of $F$. In particular, $\overline{\Theta}\in\mathrm{cPC}^{\widehat{F}}_{G}(k)$. Let $L^{\prime}$ be an unramified extension of $L$ with residue field $k^{\prime}$ and let $\mathcal{O}^{\prime}$ be the ring of integers in $L$. For $A$ in $\mathfrak{A}_{\mathcal{O}^{\prime}}$ we let $D^{\square,\widehat{F}}_{\overline{\rho}}(A)$ be the set of continuous representations $\rho:\widehat{F}\rightarrow G(A)$ congruent to $\overline{\rho}$ modulo $\mathfrak{m}_{A}$. The functor $D^{\square,\widehat{F}}_{\overline{\rho}}$ is pro-represented by a local profinite $\mathcal{O}$-algebra $R^{\square,\widehat{F}}_{\overline{\rho}}$ with residue field $k$. We employ the same notational scheme with $F$ instead of $\widehat{F}$, but drop the continuity condition.
Corollary 4.3.
The restriction maps $\Theta\mapsto\Theta|_{F}$, $\rho\mapsto\rho|_{F}$ induce isomorphisms
(7) $\mathcal{O}(\mathrm{PC}^{F}_{G})_{\mathfrak{m}}^{\wedge}\cong R^{\mathrm{ps},\widehat{F}}_{\overline{\Theta}},\quad\mathcal{O}(\operatorname{Rep}^{F}_{G})_{\mathfrak{n}}^{\wedge}\cong R^{\square,\widehat{F}}_{\overline{\rho}}$
and hence the map $\rho\mapsto\Theta_{\rho}$ induces an injection $R^{\mathrm{ps},\widehat{F}}_{\overline{\Theta}}\hookrightarrow R^{\square,\widehat{F}}_{\overline{\rho}}$.
Proof.
Any $A\in\mathfrak{A}_{\mathcal{O}^{\prime}}$ is a finite ring, and hence $G(A)$ is a finite group. Thus the restriction map $D^{\square,\widehat{F}}_{\overline{\rho}}(A)\rightarrow D^{\square,F}_{\overline{\rho}}(A)$, $\rho\mapsto\rho|_{F}$ is bijective, which implies that the induced map $R^{\square,F}_{\overline{\rho}}\rightarrow R^{\square,\widehat{F}}_{\overline{\rho}}$ is an isomorphism. Since $\mathcal{O}(\operatorname{Rep}^{F}_{G})$ is noetherian, and $\kappa(\mathfrak{n})$ is a finite field, $D^{\square,F}_{\overline{\rho}}$ is pro-represented by $\mathcal{O}(\operatorname{Rep}^{F}_{G})_{\mathfrak{n}}^{\wedge}$.
Similarly $\operatorname{Def}^{F}_{\overline{\Theta}}$ is pro-represented by $\mathcal{O}(\mathrm{PC}^{F}_{G})_{\mathfrak{m}}^{\wedge}$. Since $F$ is dense in $\widehat{F}$ the restriction $\Theta\mapsto\Theta|_{F}$ induces an injection $\operatorname{Def}^{\widehat{F}}_{\overline{\Theta}}(A)\hookrightarrow\operatorname{Def}^{F}_{\overline{\Theta}}(A)$ for all $A\in\mathfrak{A}_{\mathcal{O}}$ by [16, Lemma 3.2]. This yields a surjection $\mathcal{O}(\mathrm{PC}^{F}_{G})_{\mathfrak{m}}^{\wedge}\twoheadrightarrow R^{\mathrm{ps},\widehat{F}}_{\overline{\Theta}}$.
Since the composition
(8) $\mathcal{O}(\mathrm{PC}^{F}_{G})_{\mathfrak{m}}^{\wedge}\twoheadrightarrow R^{\mathrm{ps},\widehat{F}}_{\overline{\Theta}}\rightarrow R^{\square,\widehat{F}}_{\overline{\rho}}\cong\mathcal{O}(\operatorname{Rep}^{F}_{G})_{\mathfrak{n}}^{\wedge}$
is injective by 4.2 we obtain the assertion. ∎
Corollary 4.4.
Let $Q$ be a quotient of $\widehat{F}$ fitting in the exact sequence
(9) $0\rightarrow K(p)\rightarrow Q\rightarrow\widehat{F}/K\rightarrow 0,$
where $K\subset\widehat{F}$ is the kernel of $\overline{\rho}$ and $K(p)$ is the maximal pro-$p$ quotient of $K$. Then $R^{\mathrm{ps},\widehat{F}}_{\overline{\Theta}}=R^{\mathrm{ps},Q}_{\overline{\Theta}}$.
Proof.
If $A\in\mathfrak{A}_{\mathcal{O}}$ then the kernel of $G(A)\rightarrow G(k)$ is a finite $p$-group. Hence, if $\rho_{A}$ is a deformation of $\overline{\rho}$ to $A$ then the homomorphism $\rho_{A}:\widehat{F}\rightarrow G(A)$ factors through $Q\rightarrow G(A)$. We deduce that the surjection $R^{\square,\widehat{F}}_{\overline{\rho}}\twoheadrightarrow R^{\square,Q}_{\overline{\rho}}$ is an isomorphism. Since the natural map $R^{\mathrm{ps},\widehat{F}}_{\overline{\Theta}}\rightarrow R^{\square,\widehat{F}}_{\overline{\rho}}$ is injective by 4.3, we deduce that the composition
(10) $R^{\mathrm{ps},\widehat{F}}_{\overline{\Theta}}\twoheadrightarrow R^{\mathrm{ps},Q}_{\overline{\Theta}}\rightarrow R^{\square,Q}_{\overline{\rho}}$
is injective. Hence, the surjection $R^{\mathrm{ps},\widehat{F}}_{\overline{\Theta}}\twoheadrightarrow R^{\mathrm{ps},Q}_{\overline{\Theta}}$ is an isomorphism. ∎
Corollary 4.5.
The quotient $K/\operatorname{Ker}(\Theta^{u})$ is a pro-$p$ group.
Proof.
4.4 implies that $\widehat{F}\twoheadrightarrow\widehat{F}/\operatorname{Ker}(\Theta^{u})$ factors as $Q\twoheadrightarrow\widehat{F}/\operatorname{Ker}(\Theta^{u})$. Thus $K/\operatorname{Ker}(\Theta^{u})$ is a quotient of $K(p)$ and hence is pro-$p$. ∎
5. The general case
Let $\Gamma$ be a profinite group and let $\overline{\Theta}\in\mathrm{cPC}^{\Gamma}_{G}(k)$. Let $R^{\mathrm{ps}}:=R^{\mathrm{ps},\Gamma}_{\overline{\Theta}}$ be the universal ring representing $\operatorname{Def}_{\overline{\Theta}}$, and let $\Theta^{u}$ be the universal deformation of $\overline{\Theta}$ introduced in Section 2.3. By the reconstruction theorem [16, Theorem 3.7] there exists a $G$-semisimple representation $\overline{\rho}:\Gamma\rightarrow G(\bar{k})$, uniquely determined up to $G^{0}(\bar{k})$-conjugation, such that $\Theta_{\overline{\rho}}=\overline{\Theta}\otimes_{k}\bar{k}$. Theorem 3.8 in [16] implies that $\overline{\rho}$ is continuous. We let $K$ be the kernel of $\overline{\rho}$. Since the topology on $\bar{k}$ is discrete, $K$ is an open subgroup of $\Gamma$. Since $\overline{\rho}$ is uniquely determined up to $G^{0}(\bar{k})$-conjugation, $K$ is independent of the choice of $\overline{\rho}$. Moreover, we have inclusions $\operatorname{Ker}(\Theta^{u})\subset\operatorname{Ker}(\overline{\Theta})\subset K$ and $\operatorname{Ker}(\Theta^{u})$ is a closed normal subgroup of $\Gamma$.
Let $X\subset\Gamma$ be a set of generators for $\Gamma$ as a topological group. For a subset $Y$ of $X$ we let $F_{Y}$ be the free group generated by $Y$. The homomorphism $F_{Y}\rightarrow\Gamma$ extends to a continuous homomorphism $\widehat{F}_{Y}\rightarrow\Gamma$ from the profinite completion of $F_{Y}$. We let $\overline{\rho}_{Y}:\widehat{F}_{Y}\rightarrow\Gamma\overset{\overline{\rho}}{\longrightarrow}G(\bar{k})$, $K_{Y}$ the kernel of $\overline{\rho}_{Y}$, $\overline{\Theta}_{Y}:=\overline{\Theta}|_{\widehat{F}_{Y}}$, $R^{\mathrm{ps}}_{Y}$ the universal deformation ring representing $\operatorname{Def}_{\overline{\Theta}_{Y}}$ and $\Theta^{u}_{Y}$ the universal deformation of $\overline{\Theta}_{Y}$.
Lemma 5.1.
Let $Y$ be a subset of $X$. Then the map $\widehat{F}_{Y}\rightarrow\Gamma/\operatorname{Ker}(\Theta^{u})$ factors through $\widehat{F}_{Y}/\operatorname{Ker}(\Theta^{u}_{Y})\rightarrow\Gamma/\operatorname{Ker}(\Theta^{u})$.
Proof.
The restriction of $\Theta^{u}$ to $\widehat{F}_{Y}$ is a deformation of $\overline{\Theta}_{Y}$ to $R^{\mathrm{ps}}$. Thus there is a continuous homomorphism $R^{\mathrm{ps}}_{Y}\rightarrow R^{\mathrm{ps}}$ such that
(11) $\Theta^{u}|_{\widehat{F}_{Y}}=\Theta^{u}_{Y}\otimes_{R^{\mathrm{ps}}_{Y}}R^{\mathrm{ps}}.$
This implies the assertion. ∎
Lemma 5.2.
Let $N$ be an open normal subgroup of $K$. Then there is a finite subset $Y$ of $X$ such that $K_{Y}\rightarrow K/N$ is surjective.
Proof.
We are free to replace $N$ by an open subgroup, which is normal in $K$. Since $K$ is open in $\Gamma$ and hence of finite index, we may assume that $N$ is normal in $\Gamma$. Since $\Gamma/N$ is a finite group there is a finite subset $Y$ of $X$ such that $\widehat{F}_{Y}\rightarrow\Gamma/N$ is surjective. The preimage of $K/N$ is equal to $K_{Y}$, and hence the assertion follows. ∎
Let $Q$ be a quotient of $\Gamma$ fitting in the exact sequence
(12) $0\rightarrow K(p)\rightarrow Q\rightarrow\Gamma/K\rightarrow 0,$
where $K(p)$ is the maximal pro-$p$ quotient of $K$.
Lemma 5.3.
Let $\overline{\rho}_{1}:\Gamma\rightarrow G(\bar{k})$ be a continuous representation with $\Theta_{\overline{\rho}_{1}}=\Theta_{\overline{\rho}}$. Let $K_{1}=\operatorname{Ker}(\overline{\rho}_{1})$ and let $Q_{1}$ be the quotient of $\Gamma$ defined by (12) with $K_{1}$ instead of $K$. Then $Q=Q_{1}$.
Proof.
It follows from the reconstruction theorem that $\overline{\rho}_{1}$ and $\overline{\rho}$ have the same $G$-semisimplification. Since $\overline{\rho}$ is $G$-semisimple, after replacing $\overline{\rho}$ by a conjugate with an element of $G^{0}(\bar{k})$ we may assume that there is a parabolic subgroup $P$ of $G$ with unipotent radical $U$ and Levi $L$ such that $\overline{\rho}_{1}$ takes values in $P(\bar{k})$, $\overline{\rho}$ takes values in $L(\bar{k})$ and $\overline{\rho}(\gamma)=\overline{\rho}_{1}(\gamma)\pmod{U(\bar{k})}$ for all $\gamma\in\Gamma$. Since $U(k^{\prime})$ is a finite $p$-group for all finite extensions $k^{\prime}$ of $k$, we obtain a surjection $\Gamma/K_{1}\twoheadrightarrow\Gamma/K$ such that the kernel $K/K_{1}$ is a finite $p$-group. Thus $\Gamma/K$ is a quotient of $Q_{1}$ and $\Gamma/K_{1}$ is a quotient of $Q$ and the kernels in both cases are pro-$p$. Hence, $Q_{1}$ is a quotient of $Q$ and $Q$ is the quotient of $Q_{1}$, which implies that $Q=Q_{1}$. ∎
Proposition 5.4.
Let $N$ be a closed normal subgroup of $\Gamma$ contained in $K$. If for every finite subset $Y\subset X$ the map $K_{Y}\rightarrow\Gamma/N$ factors through the maximal pro-$p$ quotient $K_{Y}(p)$, then $K/N$ is pro-$p$. In particular, $\Gamma\twoheadrightarrow\Gamma/N$ factors through $Q\twoheadrightarrow\Gamma/N$.
Proof.
Let $N^{\prime}$ be an open normal subgroup of $K$ containing $N$. 5.2 implies that there is a finite subset $Y$ of $X$ such that $K_{Y}\rightarrow K/N^{\prime}$ is surjective. Our assumptions imply that this map factors as $K_{Y}(p)\rightarrow K/N^{\prime}$. Thus the order of $K/N^{\prime}$ is a power of $p$, which implies that $K/N$ is pro-$p$. Hence, $K\rightarrow K/N$ factors as $K(p)\rightarrow K/N$, which implies the last assertion. ∎
Corollary 5.5.
The group $K/\operatorname{Ker}(\Theta^{u})$ is pro-$p$.
Proof.
We will deduce the assertion from 5.4 applied with $N=\operatorname{Ker}(\Theta^{u})$. The map $K_{Y}\rightarrow K/\operatorname{Ker}(\Theta^{u})$ factors through $K_{Y}/\operatorname{Ker}(\Theta^{u}_{Y})$ by 5.1. Since $K_{Y}/\operatorname{Ker}(\Theta^{u}_{Y})$ is pro-$p$ by 4.5 the map $K_{Y}\rightarrow K/\operatorname{Ker}(\Theta^{u})$ factors as $K_{Y}(p)\rightarrow K/\operatorname{Ker}(\Theta^{u})$. The assertion follows from 5.4. ∎
Corollary 5.6.
$R^{\mathrm{ps},\Gamma}_{\overline{\Theta}}=R^{\mathrm{ps},Q}_{\overline{\Theta}}$.
Proof.
5.5 implies that the surjection $\Gamma\twoheadrightarrow\Gamma/\operatorname{Ker}(\Theta^{u})$ factors through $Q\twoheadrightarrow\Gamma/\operatorname{Ker}(\Theta^{u})$, which implies the assertion. ∎
6. Mazur’s finiteness condition
A profinite group $\Gamma$ satisfies Mazur’s condition $\Phi_{p}$ if $\operatorname{Hom}^{\mathrm{cont}}(\Gamma^{\prime},\mathbb{F}_{p})$ is a finite dimensional $\mathbb{F}_{p}$-vector space for every open subgroup $\Gamma^{\prime}$ of $\Gamma$. We keep the notation of the previous section.
Lemma 6.1.
If $\Gamma$ satisfies Mazur’s finiteness condition at $p$ then its quotient $Q$ in 5.6 is a finitely generated profinite group.
Proof.
Since
(13) $\operatorname{Hom}^{\mathrm{cont}}(K(p),\mathbb{F}_{p})\cong\operatorname{Hom}^{\mathrm{cont}}(K,\mathbb{F}_{p})$
we deduce that $\operatorname{Hom}^{\mathrm{cont}}(K(p),\mathbb{F}_{p})$ is a finite dimensional $\mathbb{F}_{p}$-vector space. This implies that $K(p)$ is finitely generated, [11, Lemma 2.1]. The assertion follows as $\Gamma/K$ is a finite group. ∎
5.6 and 6.1 imply that $R^{\mathrm{ps},\Gamma}_{\overline{\Theta}}=R^{\mathrm{ps},Q}_{\overline{\Theta}}$, where $Q$ is a finitely generated quotient of $\Gamma$. We may use this to transfer the results of [16], [14] and [15] proved for finitely generated profinite groups to the groups satisfying Mazur’s finiteness condition at $p$.
Corollary 6.2.
If $\Gamma$ satisfies Mazur’s condition $\Phi_{p}$ then $R^{\mathrm{ps},\Gamma}_{\overline{\Theta}}$ is noetherian.
Proof.
The assertion follows from the corresponding result for finitely generated profinite groups [16, Theorem 5.7]. ∎
6.1. The moduli space of condensed representations
Let $R$ be a profinite noetherian $\mathcal{O}$-algebra. Then any finitely generated $R$-module $M$ carries a unique Hausdorff topology, which makes it into a topological $R$-module. Let $A$ be an $R$-algebra. A representation $\rho:\Gamma\rightarrow G(A)$ is called $R$-condensed if there exists a closed immersion of $\mathcal{O}$-schemes $\tau:G\hookrightarrow\mathbb{A}^{n}$ and a finitely generated $R$-submodule $M$ of $\mathbb{A}^{n}(A)=A^{n}$ such that $\tau(\rho(\Gamma))\subset M$ and the map $\tau\circ\rho:\Gamma\rightarrow M$ is continuous for the canonical topology on $M$ as a finitely generated $R$-module. We check in [14, Lemma 4.2] that if $\rho$ is $R$-condensed with respect to one closed immersion $\tau$ then the same holds for all closed immersions. Moreover, in [14, Section 4.2] we show that $\rho$ is $R$-condensed if and only if there exists a homomorphism of condensed groups $\underline{\Gamma}\to G(A_{\operatorname{disc}}\otimes_{R_{\operatorname{disc}}}\underline{R})$ which recovers $\rho$ after evaluation at a point. Here $\underline{-}$ denotes the functor from topological spaces to condensed sets and $(-)_{\operatorname{disc}}$ takes a set to a discrete condensed set. This interpretation in terms of condensed mathematics of Clausen and Scholze [17] motivates the terminology. Informally, $R$-condensed should be thought of as a continuity condition on $\rho$, even if there is no topology on the $R$-algebra AA.
If $\Gamma$ satisfies Mazur’s condition $\Phi_{p}$, then $R^{\mathrm{ps}}_{\overline{\Theta}}$ is noetherian by 6.2. Thus we may apply the discussion above with $R=R^{\mathrm{ps}}_{\overline{\Theta}}$ and for an $R^{\mathrm{ps}}_{\overline{\Theta}}$-algebra $A$ we let $X^{\mathrm{gen}}_{\overline{\Theta}}(A)$ be the set of $R^{\mathrm{ps}}_{\overline{\Theta}}$-condensed representations $\rho:\Gamma\rightarrow G(A)$ such that $\Theta_{\rho}=\Theta^{u}\otimes_{R^{\mathrm{ps}}_{\overline{\Theta}}}A$.
Let $Q$ be the quotient of $\Gamma$ defined in (12). We will add a superscript to indicate which group we are working with. Since $R^{\mathrm{ps},\Gamma}_{\overline{\Theta}}=R^{\mathrm{ps},Q}_{\overline{\Theta}}$ by 5.6, we have an inclusion $X^{\mathrm{gen},Q}_{\overline{\Theta}}(A)\subset X^{\mathrm{gen},\Gamma}_{\overline{\Theta}}(A)$.
Proposition 6.3.
If $\Gamma$ satisfies Mazur’s condition $\Phi_{p}$ then $X^{\mathrm{gen},Q}_{\overline{\Theta}}(A)=X^{\mathrm{gen},\Gamma}_{\overline{\Theta}}(A)$.
Proof.
We choose a closed immersion of $\mathcal{O}$-group schemes $\tau:G\hookrightarrow\operatorname{GL}_{d}$. Composition with $\tau$ induces injections $X^{\mathrm{gen},\Gamma}_{\overline{\Theta}}(A)\hookrightarrow X^{\mathrm{gen},\Gamma}_{\tau\circ\overline{\Theta}}(A)$ and $X^{\mathrm{gen},Q}_{\overline{\Theta}}(A)\hookrightarrow X^{\mathrm{gen},Q}_{\tau\circ\overline{\Theta}}(A)$ by [14, Proposition 3.8]. Thus we may assume that $G=\operatorname{GL}_{d}$. In this case the assertion follows from [14, Lemma 5.5]. ∎
The proposition allows us to transfer the results proved for finitely generated profinite groups to profinite groups satisfying Mazur’s condition $\Phi_{p}$. In the corollaries below we cite the reference for the statement, when $\Gamma$ is finitely generated.
Corollary 6.4.
If $\Gamma$ satisfies Mazur’s condition $\Phi_{p}$ then the functor $X^{\mathrm{gen}}_{\overline{\Theta}}:R^{\mathrm{ps}}_{\overline{\Theta}}\rightarrow\operatorname{Set}$ is represented by a finite type $R^{\mathrm{ps}}_{\overline{\Theta}}$-algebra $A^{\mathrm{gen}}_{\overline{\Theta}}$.
Proof.
[14, Proposition 8.3]. ∎
Let $X^{\mathrm{ps}}_{\overline{\Theta}}:=\operatorname{Spec}R^{\mathrm{ps}}_{\overline{\Theta}}$. The conjugation action of $G^{0}(A)$ on $X^{\mathrm{gen}}_{\overline{\Theta}}(A)$ induces an action of $G^{0}$ on $X^{\mathrm{gen}}_{\overline{\Theta}}$. Mapping $\rho\in X^{\mathrm{gen}}_{\overline{\Theta}}(A)$ to its $G$-pseudocharacter induces a $G^{0}$-equivariant morphism $X^{\mathrm{gen}}_{\overline{\Theta}}\rightarrow X^{\mathrm{ps}}_{\overline{\Theta}}$ for the trivial action on the target. Hence the morphism factors through the GIT-quotient $X^{\mathrm{gen}}_{\overline{\Theta}}\sslash G^{0}\rightarrow X^{\mathrm{ps}}_{\overline{\Theta}}$.
Corollary 6.5.
If $\Gamma$ satisfies Mazur’s condition $\Phi_{p}$ then the natural map
$X^{\mathrm{gen}}_{\overline{\Theta}}\sslash G^{0}\rightarrow X^{\mathrm{ps}}_{\overline{\Theta}}$
is a finite adequate homeomorphism in the sense of [1, Definition 3.3.1].
Proof.
[15, Theorem 4.12]. ∎
7. Completions at maximal ideals and deformation problems
Let $\kappa$ be a local field that is an $\mathcal{O}$-algebra. Let $\mathcal{O}_{\kappa}$ be its ring of integers and let $k^{\prime}$ be the residue field. We always equip $\kappa$ with the topology induced by the valuation. Let $\mathcal{O}^{\prime}=\mathcal{O}\otimes_{W(k)}W(k^{\prime})$, where $W(k)$ is the Witt ring of $k$.
In this section we will study deformation theory of $\Theta\in\mathrm{cPC}^{\Gamma}_{G}(\kappa)$, where $\Gamma$ is an arbitrary profinite group and $G$ is a generalised reductive group scheme over $\mathcal{O}$. If $\mathrm{char}(\kappa)=0$ then we let $\Lambda=\kappa$ equipped with its natural topology and we let $\Lambda_{0}=\mathcal{O}_{\kappa}$. If $\mathrm{char}(\kappa)=p$ then [^5] The study of deformation problem in this setting is motivated by the work of Böckle–Juschka [5, Section 4.7]. we let $\Lambda=\mathcal{O}\otimes_{W(k)}C$, where $C$ is a Cohen ring of $\kappa$, which means that $C$ is a complete DVR with uniformiser $p$ and residue field $\kappa$. If we choose an isomorphism $k^{\prime}\llbracket t\rrbracket\cong\mathcal{O}_{\kappa}$ of local $k^{\prime}$-algebras then $C$ can be chosen to be the $p$-adic completion of $W(k^{\prime})\llbracket t\rrbracket[1/t]$. There is a natural topology on $\Lambda$, such that the quotient topology on $\kappa$ induced via $\Lambda\twoheadrightarrow\kappa$ coincides with the valuation topology on $\kappa$. The topology on $\Lambda$ is Hausdorff, see [4, Section 3.5] for more details. We let $\Lambda_{0}=\mathcal{O}^{\prime}\llbracket t\rrbracket$. In both cases we let $\mathfrak{A}_{\Lambda}$ be the category of local artinian $\Lambda$-algebras with residue field $\kappa$. Any $A\in\mathfrak{A}_{\Lambda}$ is a finitely generated $\Lambda$-module, and hence carries a unique Hausdorff topology, which makes $A$ into a topological $\Lambda$-module.
For $A\in\mathfrak{A}_{\Lambda}$ we let $\operatorname{Def}_{\Theta}(A)$ be the set of $\widetilde{\Theta}\in\mathrm{cPC}^{\Gamma}_{G}(A)$ such that $\widetilde{\Theta}\otimes_{A}\kappa=\Theta$. It follows [^6] If $\mathrm{char}(\kappa)=p$ then $\Lambda$ is taken to be equal to $\kappa$ in [16]. However, the argument used in the proof of [16, Theorem 5.4] goes through in our more general setting. from [16, Theorem 5.4] that the functor $\operatorname{Def}_{\Theta}:\mathfrak{A}_{\Lambda}\rightarrow\operatorname{Set}$ is pro-represented by a Hausdorff topological local $\Lambda$-algebra $R^{\mathrm{ps}}_{\Theta}$, which is a projective limit of rings in $\mathfrak{A}_{\Lambda}$. We let $\widetilde{\Theta}^{u}\in\mathrm{cPC}^{\Gamma}_{G}(R^{\mathrm{ps}}_{\Theta})$ be the the universal deformation of $\Theta$.
Lemma 7.1.
If $\Theta\in\mathrm{cPC}^{\Gamma}_{G}(\kappa)$ then $\Theta$ takes values in $\mathcal{O}_{\kappa}$.
Proof.
By the reconstruction theorem [16, Theorem 3.7] there exists a continuous representation $\rho:\Gamma\rightarrow G(\overline{\kappa})$, such that $\Theta_{\rho}=\overline{\Theta}\otimes_{\kappa}\overline{\kappa}$. If $\kappa^{\prime}$ is a finite extension of $\kappa$ contained in $\overline{\kappa}$ then $G(\kappa^{\prime})$ is a closed subgroup of $G(\overline{\kappa})$ and we let $\Gamma_{\kappa^{\prime}}$ be its preimage in $\Gamma$. Then $\Gamma_{\kappa^{\prime}}$ is a closed and hence profinite subgroup of $\Gamma$. It follows from [7, Lemma 4.3] that there exists a finite extension $\kappa^{\prime\prime}$ of $\kappa^{\prime}$ contained in $\overline{\kappa}$ and $g\in G^{0}(\kappa^{\prime\prime})$ such that $g\rho(\Gamma_{\kappa^{\prime}})g^{-1}\subset G(\mathcal{O}_{\kappa^{\prime\prime}})$. This implies that $\Theta|_{\Gamma_{\kappa^{\prime}}}$ takes values in $\mathcal{O}_{\kappa^{\prime\prime}}$. Since $\Gamma=\bigcup_{\kappa^{\prime}}\Gamma_{\kappa^{\prime}}$ we deduce that $\Theta$ takes values in $\mathcal{O}_{\overline{\kappa}}\cap\kappa=\mathcal{O}_{\kappa}$. ∎
Let $\overline{\Theta}\in\mathrm{cPC}^{\Gamma}_{G}(k^{\prime})$ be $\Theta\otimes_{\mathcal{O}_{\kappa}}k^{\prime}$. Let $R^{\mathrm{ps}}_{\overline{\Theta}}$ be the deformation ring representing $\operatorname{Def}_{\overline{\Theta}}:\mathfrak{A}_{\mathcal{O}^{\prime}}\rightarrow\operatorname{Set}$ and let $\Theta^{u}\in\mathrm{cPC}^{\Gamma}_{G}(R^{\mathrm{ps}}_{\overline{\Theta}})$ be the universal deformation of $\overline{\Theta}$. Since $\Theta$ is a deformation of $\overline{\Theta}$ to $\mathcal{O}_{\kappa}$ we obtain a homomorphism of local $\mathcal{O}^{\prime}$-algebras $\varphi:R^{\mathrm{ps}}_{\overline{\Theta}}\rightarrow\mathcal{O}_{\kappa}$. The natural map $\Lambda\otimes_{\mathcal{O}^{\prime}}R^{\mathrm{ps}}_{\overline{\Theta}}\rightarrow\kappa$ of $\Lambda$-algebras is surjective and we denote its kernel by $\mathfrak{q}$.
Proposition 7.2.
If $\Gamma$ is finitely generated then there is a natural isomorphism of local $\Lambda$-algebras between $R^{\mathrm{ps}}_{\Theta}$ and the $\mathfrak{q}$-adic completion of $\Lambda\otimes_{\mathcal{O}^{\prime}}R^{\mathrm{ps}}_{\overline{\Theta}}$. In particular, $R^{\mathrm{ps}}_{\Theta}$ is noetherian and $\widetilde{\Theta}^{u}=\Theta^{u}\otimes_{R^{\mathrm{ps}}_{\overline{\Theta}}}R^{\mathrm{ps}}_{\Theta}$.
Proof.
Let $B$ be the $\mathfrak{q}$-adic completion of $\Lambda\otimes_{\mathcal{O}^{\prime}}R^{\mathrm{ps}}_{\overline{\Theta}}$. Then $B$ is a complete local noetherian $\Lambda$-algebra with residue field $\kappa$ by [5, Lemma 3.3.5], [4, Lemma 3.36]. Thus $B/\mathfrak{m}_{B}^{n}\in\mathfrak{A}_{\Lambda}$ for all $n\geq 1$ and $\widetilde{\Theta}_{B}:=\Theta^{u}\otimes_{R^{\mathrm{ps}}_{\overline{\Theta}}}B$ is a deformation of $\Theta$ to $B$.
The ring $B$ pro-represents a subfunctor $\operatorname{Def}^{\prime}_{\Theta}$ of $\operatorname{Def}_{\Theta}$, such that if $A\in\mathfrak{A}_{\Lambda}$, then $\widetilde{\Theta}\in\operatorname{Def}_{\Theta}(A)$ lies in $\operatorname{Def}^{\prime}_{\Theta}(A)$ if and only if $\widetilde{\Theta}$ takes values in a $\Lambda_{0}$-subalgebra $A_{0}$ of $A$, which is finite as a $\Lambda_{0}$-module. Indeed, in this case $\widetilde{\Theta}$ is a deformation of $\overline{\Theta}$ to $A_{0}$, and since $A_{0}$ is a limit of rings in $\mathfrak{A}_{\mathcal{O}^{\prime}}$ we obtain a morphism of local $\mathcal{O}^{\prime}$-algebras $R^{\mathrm{ps}}_{\overline{\Theta}}\rightarrow A_{0}$, which induces a morphism $B\rightarrow A$, such that $\widetilde{\Theta}=\widetilde{\Theta}_{B}\otimes_{B}A$. Moreover, $\widetilde{\Theta}_{B}\otimes_{B}B/\mathfrak{m}_{B}^{n}$ takes values in the image of $\Lambda_{0}\otimes_{\mathcal{O}^{\prime}}R^{\mathrm{ps}}_{\overline{\Theta}}$ in $B/\mathfrak{m}_{B}^{n}$, which is a finite $\Lambda_{0}$-module.
Since $\Gamma$ is finitely generated as a topological group, we may choose a dense discrete finitely generated subgroup $\Sigma$ of $\Gamma$. Since $\Sigma$ is dense the restriction map induces an injection $\operatorname{Def}_{\Theta}(A)\subset\mathrm{PC}^{\Sigma}_{G}(A)$. Let $A_{\infty}$ be the $\Lambda_{0}$-subalgebra of $A$ equal to the preimage of $\mathcal{O}_{\kappa}$ under $A\twoheadrightarrow\kappa$. Since $\mathfrak{m}_{A}$ nilpotent and a finite $\Lambda$-module, $A_{\infty}$ is a union of $\Lambda_{0}$-subalgebras $A_{n}$ of $A$, $n\geq 1$, such that each $A_{n}$ is a finite $\Lambda_{0}$-module and the image of $A_{n}\rightarrow\kappa$ is equal to $\mathcal{O}_{\kappa}$. 7.1 implies that the restriction to $\Sigma$ identifies $\operatorname{Def}_{\Theta}(A)$ with a subset of $\mathrm{PC}^{\Sigma}_{G}(A_{\infty})$. Since $\Sigma$ is finitely generated, $\mathrm{PC}^{\Sigma}_{G}$ is represented by a finite type $\mathcal{O}$-algebra by [16, Proposition 3.21]. Hence, $\mathrm{PC}^{\Sigma}_{G}(A_{\infty})$ is the union of $\mathrm{PC}^{\Sigma}_{G}(A_{n})$ for $n\geq 1$. This implies that $\operatorname{Def}^{\prime}_{\Theta}(A)=\operatorname{Def}_{\Theta}(A)$, which implies the assertion. ∎
As in Section 5, let $\overline{\rho}:\Gamma\rightarrow G(\bar{k})$ be a continuous, $G$-semisimple representation with $\Theta_{\overline{\rho}}=\overline{\Theta}$. Let $K$ be the kernel of $\overline{\rho}$ and let $Q$ be the quotient of $\Gamma$ defined in (12).
Corollary 7.3.
If $\Gamma$ is finitely generated then the group $K/\operatorname{Ker}(\widetilde{\Theta}^{u})$ is pro-$p$ and $R^{\mathrm{ps},\Gamma}_{\Theta}=R^{\mathrm{ps},Q}_{\Theta}$.
Proof.
It follows from 7.2 that $\operatorname{Ker}(\Theta^{u})\subset\operatorname{Ker}(\widetilde{\Theta}^{u})$. Since $K/\operatorname{Ker}(\Theta^{u})$ is pro-$p$ by 5.5, $K/\operatorname{Ker}(\widetilde{\Theta}^{u})$ is also pro-pp. Thus $\Gamma/\operatorname{Ker}(\widetilde{\Theta}^{u})$ is a quotient of $Q$, which implies the equality of deformation rings. ∎
Let $X$ be a set of generators of $\Gamma$. For every subset $Y$ of $X$ the notation $K_{Y}$, $\overline{\Theta}_{Y}$, $\widehat{F}_{Y}$, $F_{Y}$ has the same meaning as in the previous section. We let $\Theta_{Y}$ be the restriction of $\Theta$ to $\widehat{F}_{Y}$, $R^{\mathrm{ps}}_{\Theta_{Y}}$ the universal deformation ring of $\Theta_{Y}$ and let $\widetilde{\Theta}^{u}_{Y}\in\mathrm{cPC}^{\widehat{F}_{Y}}_{G}(R^{\mathrm{ps}}_{\Theta_{Y}})$ be the universal deformation of $\Theta_{Y}$.
Lemma 7.4.
The map $\widehat{F}_{Y}\rightarrow\Gamma/\operatorname{Ker}(\widetilde{\Theta}^{u})$ factors as $\widehat{F}_{Y}/\operatorname{Ker}(\widetilde{\Theta}^{u}_{Y})\rightarrow\Gamma/\operatorname{Ker}(\widetilde{\Theta}^{u})$.
Proof.
The same proof as 5.1. ∎
Proposition 7.5.
The group $K/\operatorname{Ker}(\widetilde{\Theta}^{u})$ is pro-$p$ and $R^{\mathrm{ps},\Gamma}_{\Theta}=R^{\mathrm{ps},Q}_{\Theta}$.
Proof.
Since $K_{Y}/\operatorname{Ker}(\widetilde{\Theta}^{u}_{Y})$ is pro-$p$ by 7.3, the assertion follows from 5.4 applied with $N=\operatorname{Ker}(\widetilde{\Theta}^{u})$. 7.4 implies that the assumptions in 5.4 are satisfied. ∎
Corollary 7.6.
If $\Gamma$ satisfies Mazur’s condition $\Phi_{p}$ then $R^{\mathrm{ps},\Gamma}_{\Theta}$ is naturally isomorphic to the $\mathfrak{q}$-adic completion of $\Lambda\otimes_{\mathcal{O}^{\prime}}R^{\mathrm{ps},\Gamma}_{\overline{\Theta}}$. In particular, $R^{\mathrm{ps},\Gamma}_{\Theta}$ is noetherian.
Proof.
We have $R^{\mathrm{ps},\Gamma}_{\overline{\Theta}}=R^{\mathrm{ps},Q}_{\overline{\Theta}}$ by 5.6 and $R^{\mathrm{ps},\Gamma}_{\Theta}=R^{\mathrm{ps},Q}_{\Theta}$ by 7.5. Since $\Gamma$ satisfies $\Phi_{p}$, $Q$ is finitely generated and the assertion follows from 7.2. ∎
8. The rigid analytic space of $G$-pseudocharacters
Let $\mathrm{Rig}_{L}$ be the category of rigid analytic spaces over $L$. If $Y\in\mathrm{Rig}_{L}$ then its ring of global sections $\mathcal{O}(Y)$ carries a natural Hausdorff topology. If $Y=\operatorname{Sp}(A)$ is affinoid then the topology on $A$ is the Banach space topology. Let $\tilde{X}_{G}:\mathrm{Rig}_{L}^{\mathrm{op}}\rightarrow\operatorname{Set}$ be the functor that associates to every rigid analytic space $Y\in\mathrm{Rig}_{L}$ the set of continuous $G$-pseudocharacters $\mathrm{cPC}^{\Gamma}_{G}(\mathcal{O}(Y))$. We show that if $\Gamma$ satisfies Mazur’s condition $\Phi_{p}$, then $\tilde{X}_{G}$ is represented by a rigid analytic space.
Proposition 8.1.
Let $\Theta\in\mathrm{cPC}^{\Gamma}_{G}(A)$, where $A$ is an affinoid $L$-algebra. Then $\Gamma/\operatorname{Ker}(\Theta)$ is an extension of a finite group by a pro-$p$ group.
Proof.
Let $\mathfrak{p}_{1},\ldots,\mathfrak{p}_{r}$ be the associated primes of $A$. For each $i$ we choose a maximal ideal $\mathfrak{m}_{i}$ containing $\mathfrak{p}_{i}$. The natural map $A\rightarrow\prod_{i=1}^{r}A_{\mathfrak{m}_{i}}$ is injective since the map $A\rightarrow\prod_{i=1}^{r}A_{\mathfrak{p}_{i}}$ is injective. Thus the map $A\rightarrow\prod_{i=1}^{r}\hat{A}_{\mathfrak{m}_{i}}$ is injective.
Let $\kappa_{i}$ be the residue field of $\mathfrak{m}_{i}$. Since $A$ is an affinoid algebra, $\kappa_{i}$ is a finite extension of $\mathbb{Q}_{p}$. Let $\Theta_{i}=\Theta\otimes_{A}\kappa_{i}$ then $\Theta_{i}\in\mathrm{cPC}^{\Gamma}_{G}(\kappa_{i})$ and $\widetilde{\Theta}_{i}=\Theta\otimes_{A}\hat{A}_{\mathfrak{m}_{i}}$ is a deformation of $\Theta_{i}$ to $\hat{A}_{\mathfrak{m}_{i}}$. 7.5 implies that $\Gamma/\operatorname{Ker}(\widetilde{\Theta}_{i})$ is an extension of a finite group by a pro-$p$ group. Hence the same applies to $\Gamma/\bigcap_{i=1}^{r}\operatorname{Ker}(\widetilde{\Theta}_{i})$. Since $A$ injects into $\prod_{i=1}^{r}\hat{A}_{\mathfrak{m}_{i}}$ we conclude that $\Gamma/\operatorname{Ker}(\Theta)$ is a quotient of $\Gamma/\bigcap_{i=1}^{r}\operatorname{Ker}(\widetilde{\Theta}_{i})$ and the assertion follows. ∎
Corollary 8.2.
Let $\Theta\in\mathrm{cPC}^{\Gamma}_{G}(A)$, where $A$ is an affinoid $L$-algebra. If $\Gamma$ satisfies Mazur’s condition $\Phi_{p}$, then there exists an $\mathcal{O}$-subalgebra $A_{0}\subseteq A$ which is a quotient of $\mathcal{O}\langle x_{1},\dots,x_{t}\rangle$ and $A_{0}[1/p]=A$, such that $\Theta$ takes values in $A_{0}$.
Proof.
Our assumption on $\Gamma$ and 8.1 imply that $\Gamma/\operatorname{Ker}(\Theta)$ is finitely generated. The assertion follows from [16, Lemma 6.16 (2)]. ∎
Let $|\mathrm{PC}^{\Gamma}_{G}|\subset\mathrm{PC}^{\Gamma}_{G}$ be the subset of closed points $z$ with finite residue field $k_{z}$, such that the canonical $G$-pseudocharacter $\Theta_{z}\in\mathrm{PC}^{\Gamma}_{G}(k_{z})$ attached to $z$ is continuous for the discrete topology on $k_{z}$. Let $\mathcal{O}_{z}=\mathcal{O}\otimes_{W(k)}W(k_{z})$ and let $R^{\mathrm{ps}}_{\Theta_{z}}$ be the deformation ring representing $\operatorname{Def}_{\Theta_{z}}:\mathfrak{A}_{\mathcal{O}_{z}}\rightarrow\operatorname{Set}$ and let $\Theta^{u}_{z}\in\mathrm{cPC}^{\Gamma}_{G}(R^{\mathrm{ps}}_{\Theta_{z}})$ be the universal deformation of $\Theta_{z}$. If $\Gamma$ satisfies Mazur’s condition $\Phi_{p}$ then $R^{\mathrm{ps}}_{\Theta_{z}}$ is a complete local noetherian $\mathcal{O}_{z}$-algebra with residue field $k_{z}$ by 6.2, and we let $(\operatorname{Spf}R^{\mathrm{ps}}_{\Theta_{z}})^{\mathrm{rig}}$ be the Berthelot rigid generic fibre of the formal scheme $\operatorname{Spf}R^{\mathrm{ps}}_{\Theta_{z}}$, [9, Section 7].
Theorem 8.3.
If $\Gamma$ satisfies Mazur’s condition $\Phi_{p}$, then the functor $\tilde{X}_{G}$ is representable by a quasi-Stein space $X_{G}:=\coprod_{z\in|\mathrm{PC}^{\Gamma}_{G}|}(\operatorname{Spf}R^{\mathrm{ps}}_{\Theta_{z}})^{\mathrm{rig}}$.
Proof.
Let $(R^{\mathrm{ps}}_{\Theta_{z}})^{\mathrm{rig}}$ be the ring of global sections of $(\operatorname{Spf}R^{\mathrm{ps}}_{\Theta_{z}})^{\mathrm{rig}}$. Then $\mathcal{O}(X_{G})=\prod_{z}(R^{\mathrm{ps}}_{\Theta_{z}})^{\mathrm{rig}}$ and $\Theta^{\mathrm{univ}}:=\prod_{z}\Theta^{u}_{z}\otimes_{R^{\mathrm{ps}}_{\Theta_{z}}}(R^{\mathrm{ps}}_{\Theta_{z}})^{\mathrm{rig}}\in\tilde{X}_{G}(X_{G})$. A morphism of rigid analytic spaces $Y\rightarrow X_{G}$ induces a map of topological $L$-algebras $\mathcal{O}(X_{G})\rightarrow\mathcal{O}(Y)$ to which we may associate $\Theta^{\mathrm{univ}}\otimes_{\mathcal{O}(X_{G})}\mathcal{O}(Y)\in\tilde{X}_{G}(Y)$. We claim that the map $\operatorname{Mor}_{\mathrm{Rig}_{L}}(Y,X_{G})\rightarrow\tilde{X}_{G}(Y)$ defined above is bijective. It is enough to prove the claim for $Y$ affinoid, since the general case follows by considering admissible coverings.
Let $A$ be an $L$-affinoid algebra, and let $\Theta\in\mathrm{cPC}^{\Gamma}_{G}(A)$. It follows from 8.1 that $\Theta\in\mathrm{cPC}^{Q}_{G}(A)$ for a finitely generated quotient $Q$ of $\Gamma$. Since the assertion of the theorem for finitely generated groups $\Gamma$ is proved in [16, Theorem 6.1], $\Theta$ induces a morphism of rigid analytic spaces:
(14) $\operatorname{Sp}(A)\rightarrow\coprod_{z\in|\mathrm{PC}^{Q}_{G}|}(\operatorname{Spf}R^{\mathrm{ps},Q}_{\Theta_{z}})^{\mathrm{rig}}\subseteq\coprod_{z\in|\mathrm{PC}^{\Gamma}_{G}|}(\operatorname{Spf}R^{\mathrm{ps},\Gamma}_{\Theta_{z}})^{\mathrm{rig}}$
such that $\Theta=\Theta^{\mathrm{univ}}\otimes_{\mathcal{O}(X_{G})}A$. Thus we obtain a canonical bijection between $\tilde{X}_{G}(\operatorname{Sp}(A))$ and $\operatorname{Mor}_{\mathrm{Rig}_{L}}(\operatorname{Sp}(A),X_{G})$. ∎
Corollary 8.4.
If $\Gamma$ satisfies Mazur’s condition $\Phi_{p}$ then the $\overline{L}$-points of $X_{G}$ are in canonical bijection with $G^{0}(\overline{L})$-conjugacy classes of continuous $G$-semisimple representations $\Gamma\rightarrow G(\overline{L})$.
Proof.
If $\Gamma$ is finitely generated then the assertion is proved in [16, Corollary 6.23] and the same proof carries over. By the definition of $X_{G}$, we have $X_{G}(\overline{L})=\mathrm{cPC}^{\Gamma}_{G}(\overline{L})$, where $\overline{L}$ carries the direct limit topology of its finite subextensions of $L$. The claim now follows from the reconstruction theorem [16, Theorem 3.8]. ∎
Remark 8.5.
A priori it is not at all clear that the set of $G^{0}(\overline{L})$-conjugacy classes of continuous $G$-semisimple representations $\rho:\Gamma\rightarrow G(\overline{L})$ can be parameterised by $\overline{L}$-points of a rigid analytic space. The finiteness statements proved in 6.2 and 8.3 are key inputs in the proof of this statement.